OpenStudy (anonymous):

Triangle ABC is located at A(-4, 5), B (-4, -1), and C(-1,1). A translation of the triangle is located at A’(-2, 8), B’(-2, 2), and C’(1, 4). How is the triangle translated? A. Triangle ABC is translated 3 units left and 2 units down. B. Triangle ABC is translated 2 units left and 3 units up. C. Triangle ABC is translated 2 units right and 3 units up. D. Triangle ABC is translated 3 units right and 2 units up.

OpenStudy (anonymous):

we have :  We have it in all . and if you calculate them can answer to your self .

OpenStudy (anonymous):

@brittanyburton : Understooooooooooooooooooooooooooooond ??????????????:D

OpenStudy (anonymous):

that picture makes no sense to me

OpenStudy (anonymous):

What number all of the xs and ys sum in ? for ys = +3 and Xs = +2 Right ?:)

OpenStudy (anonymous):

@brittanyburton

OpenStudy (anonymous):

i have no idea. wth is ys=+3

OpenStudy (anonymous):

Y2-y1=8-5=3 all the Y sum in +3 . Got it ????

OpenStudy (anonymous):

@brittanyburton

OpenStudy (anonymous):

no i dont get it sorrry

OpenStudy (anonymous):

If we find they sum in which number can answer right ?:)

OpenStudy (anonymous):

honestly, i just want the answer but also i dont get what you're saying at all. I should know bc I'm out of this math but im doing someones virtual high school for $

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

k ! It s not good but we had : \[DeltaY=3\] and Delta x = +2 So its C .!:)

OpenStudy (anonymous):

i was close i had B lol